# Search for neutrino emission from blazar $\gamma$-ray flares accounting for possible neutrino time delays

### v1

@@ 5 Discussion and conclusion
The two statistical tests weighing blazar-associated neutrino alerts accounting for the time lag of the neutrino arrival due to slowness of the proton energy losses in $p\gamma$ interactions and/or proton acceleration resulted in no evidence of a significant correlation between IceCat-1 alerts and the AGNs from H21+ or R24+ source lists. The low statistical significance of these associations $(\lesssim 2\sigma)$, given the small size, is in line with predictions of [60] and implies [25, 46, 49] that only a small fraction of AGNs can be associated with IceCube alerts.
@@
  • Blazars are inefficient sources of $\sim 100$ TeV neutrinos. Indeed, in Ref. [27], we estimated from an extrapolation of a leptohadronic model for the brightest blazar flare of 3C 454.3 to the whole population of Fermi-LAT FSRQs that, on long-term average, they produce $\sim\left(\rho_{p/e}/130\right)\times 0.5$% of IceCube neutrinos at energies greater than $100$ TeV, where $\rho_{p/e}$ is the ratio of the proton and electron energy densities, which is a proxy of the baryon loading factor (the ratio of the proton and $\gamma$-ray luminosities). Taking this rough estimate at face value, given $N_{\nu}=348$ IceCat-1 alerts with their median signalness of $0.41$, we obtain $0.71$ expected neutrinos from Fermi-LAT FSRQs, which naively implies $\mathcal{O}(1)$ AGN-neutrino associations, assuming that other types of Fermi-LAT AGNs contribute to the neutrino flux similarly.
  • Blazars are efficient neutrino sources at energies $\gtrsim 10$ PeV. If blazars are efficient accelerators of protons [^8] By this we mean the case when protons are accelerated close to the Hillas limit [61], when the ratio of the proton acceleration timescale to the gyroperiod at the maximum energy $\xi_{\mathrm{acc}}\lesssim 10^{3}$., then the peak of the neutrino spectral-energy distributions will be around $E_{\nu}\gtrsim 100$ PeV because the protons of similar energies in the jet frame reach the $p\gamma$ reaction threshold and begin efficiently interacting with the abundant infrared photons (see, e.g., fig. 12 in [27]). Ref. [29] also shows that this allows to decrease the proton power below the Eddington limit alleviating the issue with super-Eddington proton luminosities often required to explain $\sim 100$ TeV neutrinos.
@@
  • Even if $t^{\prime}_{\mathrm{delay}}$ is universal for all neutrino sources, its transformation from the jet frame to the observer frame as $t^{\prime}_{\mathrm{delay}}(1+z)/D$ is significantly affected by the value of $D$. Indeed, there are $174$ AGNs which are present both in the H21+ and R24+ source lists, and the mean ratio of values from H21+ and R24+ is $\langle D^{\mathrm{H}21+}/D^{\mathrm{R}24+}\rangle=1.8$ and the standard deviation of $\sigma(D^{\mathrm{H}21+}/D^{\mathrm{R}24+})=2.1$ indicating a rather large discrepancy between the values obtained with the two different methods. The differences in estimates of Doppler factors in blazar jets with different methods are well known [62], and even within the same method, at different observational epochs, the same source may have various values of $D$ [63]. Our adoption of the Gaussian weight accounting for the delay was meant to allow some room for uncertainty in $D$ but if it is as large as $\sim 100$%, this can hinder attempts to reveal a significant correlation between major $\gamma$-ray flares and arrivals of IceCube neutrinos.
@@
  • Proton acceleration and/or energy dissipation are decoupled from those of electrons. As Ref. [64] pointed out, in many leptohadronic models, the hadronic contribution is not needed to describe the observational electromagnetic data well [65, 66, 67, 68, 69, 50, 70, 71, 56, 72, 73]. This implies that potential neutrino emission even if some hadrons are present in the jet might be not causally connected to the leptonic emission because, e.g., protons are accelerated in a different region of the jet and/or most efficiently dissipate in zones different from those of electrons. This, in turn, leads to (i) neutrino flux not necessarily related to the electromagnetic flux of the source; (ii) no characteristic time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino production and electromagnetic flares.
@@
Refs. [12, 17, 19] found an indication of the association of IceCube high-energy neutrinos with radio flares (see, however, recent work [74], where such a correlation was not found), which are observed, on average, to lag behind $\gamma$-ray flares by several months in the Earth frame [26] as expected in some models [75, 76, 77]. Future studies utilising both $\gamma$-ray and radio light curves may increase the sensitivity of searches for neutrino-blazar associations. The releases of the new CAZ catalogue with $858$ Doppler factors [78] and updated IceCat-2 catalogue of IceCube alerts [79] will significantly enlarge both the sizes of the ensembles of sources and neutrino alerts allowing to improve the sensitivity of the search for potentially delayed neutrinos.
@@
Other projected or constructed neutrino telescopes, such as the Pacific Ocean Neutrino Experiment [80], the Baikal Gigaton Volume Detector [81], and the Cubic Kilometre Neutrino Telescope [82], Tropical Deep-sea Neutrino Telescope [83], IceCube-Gen2 [84, 85], and the Radio Neutrino Observatory in Greenland [86] will help to provide the definitive answer to the question about the total contribution of the AGNs to the astrophysical neutrino flux and a role of the possible neutrino delay with respect to the electromagnetic emission.
@@
In conclusion, the two cross-matching tests between IceCat-1 neutrino alerts and AGNs from the MOJAVE-based source sample H21+ [37] and CGRaBS-based source sample R24+ [34] with scanning over the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino arrival time in IceCube and prior major $\gamma$-ray flares resulted in no significant evidence of the correlation.
@@ Acknowledgments
This research has made use of the Fermi-LAT [87, 41, 40, 32] data provided by NASA Goddard Space Flight Center, and of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the MOJAVE database, which is maintained by the MOJAVE team [88], and the data from the All-Sky Survey of Gamma-Ray Blazar Candidates CGRaBS [44]. This work has made use of the following software packages: Python [89], AstroPy [90, 91, 92], Jupyter [93, 94], matplotlib [95], NumPy [96], SciPy [97, 98], astroquery [99, 100], pandas [101, 102], snakemake [103], pyLCR [32] [^9] https://github.com/dankocevski/pyLCR, and tqdm [104]. This research has made use of NASA’s Astrophysics Data System, INSPIRE-HEP, and Pathfinder [105]. Software citation information aggregated using the Software Citation Station [106, 107] [^10] https://www.tomwagg.com/software-citation-station/.

### v2

@@ 5 Discussion and conclusion
The two statistical tests weighing blazar-associated neutrino alerts accounting for the time lag of the neutrino arrival due to slowness of the proton energy losses in $p\gamma$ interactions and/or proton acceleration resulted in no evidence of a significant correlation between IceCat-1 alerts and the AGNs from H21+ or R24+ source lists. The low statistical significance of these associations $(\lesssim 2\sigma)$, given the small size, is in line with predictions of Liodakis et al. (2022) and implies (Abbasi and others, 2023c; Kouch et al., 2025b, a) that only a small fraction of AGNs can be associated with IceCube alerts.
@@
  • Blazars are inefficient sources of $\sim 100$ TeV neutrinos. Indeed, in our recent work (Podlesnyi and Oikonomou, 2025), we estimated from an extrapolation of a leptohadronic model for the brightest blazar flare of 3C 454.3 to the whole population of Fermi-LAT FSRQs that, on long-term average, they produce $\sim\left(\rho_{p/e}/130\right)\times 0.5$% of IceCube neutrinos at energies greater than $100$ TeV, where $\rho_{p/e}$ is the ratio of the proton and electron energy densities, which is a proxy of the baryon loading factor (the ratio of the proton and $\gamma$-ray luminosities). Taking this rough estimate at face value, given $N_{\nu}=348$ IceCat-1 alerts with their median signalness of $0.41$, we obtain $0.71$ expected neutrinos from Fermi-LAT FSRQs, which naively implies $\mathcal{O}(1)$ AGN-neutrino associations, assuming that other types of Fermi-LAT AGNs contribute to the neutrino flux similarly.
  • Blazars are efficient neutrino sources at energies $\gtrsim 10$ PeV. If blazars are efficient accelerators of protons [^7] By this we mean the case when protons are accelerated close to the Hillas limit (Hillas, 1984), when the ratio of the proton acceleration timescale to the gyroperiod at the maximum energy $\xi_{\mathrm{acc}}\lesssim 10^{3}$., then the peak of the neutrino spectral-energy distributions will be around $E_{\nu}\gtrsim 100$ PeV because the protons of similar energies in the jet frame reach the $p\gamma$ reaction threshold and begin efficiently interacting with the abundant infrared photons (see, e.g., Fig. 12 by Podlesnyi and Oikonomou, 2025). Rodrigues et al. (2025) also shows that this allows to decrease the proton power below the Eddington limit alleviating the issue with super-Eddington proton luminosities often required to explain $\sim 100$ TeV neutrinos.
@@
  • Even if $t^{\prime}_{\mathrm{delay}}$ is universal for all neutrino sources, its transformation from the jet frame to the observer frame as $t^{\prime}_{\mathrm{delay}}(1+z)/D$ is significantly affected by the value of $D$. Indeed, there are $174$ AGNs which are present both in the H21+ and R24+ source lists, and the mean ratio of values from H21+ and R24+ is $\langle D^{\mathrm{H}21+}/D^{\mathrm{R}24+}\rangle=1.8$ and the standard deviation of $\sigma(D^{\mathrm{H}21+}/D^{\mathrm{R}24+})=2.1$ indicating a rather large discrepancy between the values obtained with the two different methods. The differences in estimates of Doppler factors in blazar jets with different methods are well known (Liodakis et al., 2017), and even within the same method, at different observational epochs, the same source may have various values of $D$ (Jorstad and others, 2013). Our adoption of the Gaussian weight accounting for the delay was meant to allow some room for uncertainty in $D$ but if it is as large as $\sim 100$%, this can hinder attempts to reveal a significant correlation between major $\gamma$-ray flares and arrivals of IceCube neutrinos.
@@
  • Proton acceleration and/or energy dissipation are decoupled from those of electrons. As Capel et al. (2022) pointed out, in many leptohadronic models, the hadronic contribution is not needed to describe the observational electromagnetic data well (Keivani and others, 2018; Ansoldi and others, 2018; Murase et al., 2018; Sahakyan, 2018; Gao et al., 2019; Oikonomou et al., 2019; Cerruti et al., 2019; Reimer et al., 2019; Rodrigues et al., 2019; Petropoulou and others, 2020; Oikonomou et al., 2021). This implies that potential neutrino emission even if some hadrons are present in the jet might be not causally connected to the leptonic emission because, e.g., protons are accelerated in a different region of the jet and/or most efficiently dissipate in zones different from those of electrons. This, in turn, leads to (i) neutrino flux not necessarily related to the electromagnetic flux of the source; (ii) no characteristic time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino production and electromagnetic flares.
@@
Plavin et al. (2020); Hovatta and others (2021); Kouch and others (2024) found an indication of the association of IceCube high-energy neutrinos with radio flares (see, however, recent work by Abbasi and others (2025a), where such a correlation was not found), which are observed, on average, to lag behind $\gamma$-ray flares by several months in the Earth frame (Kramarenko et al., 2021) as expected in some models (Boula and Mastichiadis, 2022; Tramacere et al., 2022; Zacharias, 2023). Future studies utilising both $\gamma$-ray and radio light curves may increase the sensitivity of searches for neutrino-blazar associations. The releases of the new CAZ catalogue with $858$ Doppler factors (Kouch and others, 2025) and updated IceCat-2 catalogue of IceCube alerts (Abbasi and others, 2025b) will significantly enlarge both the sizes of the ensembles of sources and neutrino alerts allowing to improve the sensitivity of the search for potentially delayed neutrinos.
@@
Other projected or constructed neutrino telescopes, such as the Pacific Ocean Neutrino Experiment (Agostini and others, 2020), the Baikal Gigaton Volume Detector (Avrorin and others, 2011), and the Cubic Kilometre Neutrino Telescope (Adrian-Martinez and others, 2016), Tropical Deep-sea Neutrino Telescope (Ye and others, 2023), IceCube-Gen2 (Aartsen and others, 2021; Ishihara and others, 2023), and the Radio Neutrino Observatory in Greenland (Aguilar and others, 2021) will help to provide the definitive answer to the question about the total contribution of the AGNs to the astrophysical neutrino flux and a role of the possible neutrino delay with respect to the electromagnetic emission.
@@
In conclusion, the two cross-matching tests between IceCat-1 neutrino alerts and AGNs from the MOJAVE-based source sample H21+ (Homan et al., 2021) and CGRaBS-based source sample R24+ (Rodrigues et al., 2024b) with scanning over the jet-frame time delay $t^{\prime}_{\mathrm{delay}}$ between the neutrino arrival time in IceCube and prior major $\gamma$-ray flares resulted in no significant evidence of the correlation.
@@ Acknowledgments
This research has made use of the Fermi-LAT (Atwood and others, 2009; Ajello and others, 2020; Abdollahi and others, 2020, 2023) data provided by NASA Goddard Space Flight Center, and of the NASA/IPAC Extragalactic Database, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of data from the MOJAVE database, which is maintained by the MOJAVE team (Lister et al., 2018), and the data from the All-Sky Survey of Gamma-Ray Blazar Candidates CGRaBS (Healey et al., 2008). This work has made use of the following software packages: Python (Van Rossum and Drake, 2009), AstroPy (Astropy Collaboration, 2013; Astropy Collaboration and Astropy Contributors, 2018; Astropy Collaboration, 2022), Jupyter (Perez and Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), NumPy (Harris and others, 2020), SciPy (Virtanen and others, 2020; Pauli Virtanen et al., 2021), astroquery (Astroquery collaboration and a subset of the astropy collaboration, 2019; Adam Ginsburg et al., 2024), pandas (pandas development team, 2020; McKinney, 2010), snakemake (Mölder et al., 2021), pyLCR (Abdollahi and others, 2023) [^8] https://github.com/dankocevski/pyLCR, and tqdm (Casper da Costa-Luis et al., 2023). This research has made use of NASA’s Astrophysics Data System, INSPIRE-HEP, and Pathfinder (Kartheik G. Iyer et al., 2024). Software citation information aggregated using the Software Citation Station (Wagg and Broekgaarden, 2024b, a) [^9] https://www.tomwagg.com/software-citation-station/.